# A Method to Improve the Performance of Reinforcement Learning Based on the $\mathcal{Y}$ Operator for a Class of Stochastic Differential Equation-Based Child-Mother Systems

### v3

@@ 6.1 Linear Numerical Examples
[…]
@@
It is clear that the training performance of the YORL method using the $\mathcal{Y}$ operator is outperformed by the TSRL method in this linear child-mother system and the YORL method is able to achieve higher rewards than the TSRL method at the time of training convergence. This is due to the fact that YORL incorporates the stochasticity of the child-mother system modeled by SDEs into the design of the Critic network, and therefore is outperforming the TSRL, a method that does not take into account the stochasticity in the model, in terms of the reward function. It nicely addresses the deficiencies of the current study as articulated in Problem I. As can be seen in Fig.1, when the Relu function is used for the activation function of the network, there is a certain degree of reduction in the reward function for both the YORL and TSRL methods, in consistent with what was described in Problem II. However, the value of reward function of the YORL method is still higher than that of the conventional TSRL method because the $\mathcal{Y}$ operator avoids the problem of calculating the partial differentiation of the state value function. The experimental results reveal that the network with Relu activation function has faster convergence characteristics when the hidden layer dimensions are the same. While the network with Sigmoid activation function is slower in convergence, the peak size of the reward after convergence performs better compared to other networks. Compared to Relu and Sigmoid, the network with tanh activation function has a more moderate performance. In addition, the TSRL method converges faster than YORL when the hidden layer dimension increases, but the average reward after convergence is inferior to the YORL method.
@@ 6.2 Nonlinear Numerical Examples
The simulation results for this nonlinear child-mother system are shown in Fig.2 . Similar to the linear child-mother system, the dimension of the hidden layer used in subfigure (a) is $32$, and the activation function is a Sigmoid function. It can be found that the average reward of YORL is totally higher than the existing TSRL method throughout the training process. When the hidden layer dimension is increased to $128$ and the activation function is unchanged, as shown in subfigure (b), both the TSRL method and the YORL are improved, which is a positive result of the increase of neurons in the neural network. When the training converges, it can be noticed that the average reward of the YORL method is slightly better than the TSRL method. When the dimension of the hidden layer is unchanged and the activation function is replaced with the Relu function, as shown in subfigure (c), the advantage of YORL is still obvious. Finally, in subfigure (d), the hidden layer dimension is still $32$ and the activation function is replaced with the tanh function. YORL continues to perform remarkably well. From the two numerical examples, it can be seen that the YORL method shows better control performance than the TSRL method, both in systems modeled by linear SDEs and in systems modeled by nonlinear SDEs.The YORL method solves Problem I well and gives a feasible approach in some special scenarios described in Problem II. The YORL method can provide an illuminating proposal for researchers when the problem they are studying is in the special scenarios described in Problem II.
@@ 7 CONCLUSIONS
It can be found that YORL method outperforms TSRL method in both linear child-mother systems and nonlinear child-mother systems, which is a well-suited answer to the Problem I posed in the previous section. This demonstrates the necessity of taking the stochasticity in the system into account in the design of the loss function when designing Critic networks. When the states in the child-mother system are all considered as stochastic processes represented by SDEs, the state-value function is essentially a functional of the stochastic process. The characteristic operator of the Itô diffusion process $\mathcal{A}$ transforms the problem of solving the derivative of this functional with respect to time into a problem of solving the partial derivative of this functional with respect to its independent variables. The $\mathcal{Y}$ operator proposed in this paper can transform the problem of solving the derivative of the functional with respect to time into the problem of solving the partial derivatives of the drift term function and the diffusion term function of SDEs of the stochastic process with respect to its independent variables. For solving the transformed problem of the $\mathcal{A}$ operator, it is required that the functional of the stochastic process is known and there is a strict requirement on the continuity of the functional (at least the second-order partial derivatives are required to be continuous). However, in many scenarios, the functional may not be known, and only the data for the inputs and outputs of that functional are available. In designing a Critic network for reinforcement learning, some activation functions such as Relu may not satisfy the continuity requirement, suggesting that the choice of the activation function may result in the functional of the value function that does not satisfy the conditions for the application of the $\mathcal{A}$ operator, which is a significant limitation for the design of the network. The $\mathcal{Y}$ operator proposed in this paper transforms this problem into a problem of solving the partial derivatives of the drift term function and the diffusion term function of SDEs obeyed by the random variable itself with respect to its independent variables. The continuity requirement of the pairwise value function is transformed into the continuity requirement of the drift term function and the diffusion term function of SDEs obeyed by the random variable itself. The proposed $\mathcal{Y}$ operator allows more selectivity in the design of the Critic network when the SDEs obeyed by the system state satisfy the continuity condition. The activation functions of network such as the Relu function can be chosen. Moreover, the $\mathcal{Y}$ operator can be used as a reference tool in areas such as inverse reinforcement learning, offline reinforcement learning, to provide researchers with inspiring proposals. Therefore, the proposed $\mathcal{Y}$ operator effectively addresses the shortcomings articulated in Problem II.
@@ 8.1 The Proof of Proposition 1
By denoting the dimension of $X^{(1)}_{t}$ as $d_{x^{1}}$, the dimension of $X^{(2)}_{t}$ as $d_{x^{2}}$, and so on, and the dimension of $X^{(m)}_{t}$ as $d_{x^{m}}$, there are, in essence, $\sum_{i=1}^{m}d_{x^{i}}$ independent variables in the function $\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})$. Denoted $\sum_{i=1}^{m}d_{x^{i}}$ as $d_{M}$ According to the Taylor expansion formula for multivariate functions, (34) in Proposition 1 can be written in the following form:
@@
    Case1($i\neq k$): Since $dW^{(1)}_{t}$, $dW^{(2)}_{t}$dots $dW^{(m)}_{t}$ are independent of each other, it is clear that the following equation holds
@@ 8.2 The Proof of Proposition 3
According to Theorem 1, the (36) naturally holds. And according to Kolmogorov forward equation,

### v4

@@ 6.1 Linear Numerical Examples
    Actor Network     $1\times\text{hidden dimension}\times 1$     $c_{1}$     55
    Critic Network     $1\times\text{hidden dimension}\times\text{hidden dimension}\times 1$     $c_{2}$     $0.1$
@@
It is clear that the training performance of the YORL method using the $\mathcal{Y}$ operator is outperformed by the TSRL method in this linear child-mother system and the YORL method is able to achieve higher rewards than the TSRL method at the time of training convergence. This is due to the fact that YORL incorporates the stochasticity of the child-mother system modeled by SDEs into the design of the Critic network, and therefore is outperforming the TSRL, a method that does not take into account the stochasticity in the model, in terms of the reward function. It nicely addresses the deficiencies of the current study as articulated in Problem I.
As can be seen in Fig.1, when the Relu function is used for the activation function of the network, there is a certain degree of reduction in the reward function for both the YORL and TSRL methods, in consistent with what was described in Problem II. However, the value of reward function of the YORL method is still higher than that of the conventional TSRL method because the $\mathcal{Y}$ operator avoids the problem of calculating the partial differentiation of the state value function.
The experimental results reveal that the network with Relu activation function has faster convergence characteristics when the hidden layer dimensions are the same. While the network with Sigmoid activation function is slower in convergence, the peak size of the reward after convergence performs better compared to other networks. Compared to Relu and Sigmoid, the network with tanh activation function has a more moderate performance. In addition, the TSRL method converges faster than YORL when the hidden layer dimension increases, but the average reward after convergence is inferior to the YORL method.
@@ 6.2 Nonlinear Numerical Examples
The simulation results for this nonlinear child-mother system are shown in Fig.2 . Similar to the linear child-mother system, the dimension of the hidden layer used in subfigure (a) is $32$, and the activation function is a Sigmoid function. It can be found that the average reward of YORL is totally higher than the existing TSRL method throughout the training process. When the hidden layer dimension is increased to $128$ and the activation function is unchanged, as shown in subfigure (b), both the TSRL method and the YORL are improved, which is a positive result of the increase of neurons in the neural network. When the training converges, it can be noticed that the average reward of the YORL method is slightly better than the TSRL method. When the dimension of the hidden layer is unchanged and the activation function is replaced with the Relu function, as shown in subfigure (c), the advantage of YORL is still obvious. Finally, in subfigure (d), the hidden layer dimension is still $32$ and the activation function is replaced with the tanh function. YORL continues to perform remarkably well.
From the two numerical examples, it can be seen that the YORL method shows better control performance than the TSRL method, both in systems modeled by linear SDEs and in systems modeled by nonlinear SDEs.The YORL method solves Problem I well and gives a feasible approach in some special scenarios described in Problem II. The YORL method can provide an illuminating proposal for researchers when the problem they are studying is in the special scenarios described in Problem II.
@@ 7 CONCLUSIONS
It can be found that YORL method outperforms TSRL method in both linear child-mother systems and nonlinear child-mother systems, which is a well-suited answer to the Problem I posed in the previous section. This demonstrates the necessity of taking the stochasticity in the system into account in the design of the loss function when designing Critic networks.
When the states in the child-mother system are all considered as stochastic processes represented by SDEs, the state-value function is essentially a functional of the stochastic process. The characteristic operator of the Itô diffusion process $\mathcal{A}$ transforms the problem of solving the derivative of this functional with respect to time into a problem of solving the partial derivative of this functional with respect to its independent variables. The $\mathcal{Y}$ operator proposed in this paper can transform the problem of solving the derivative of the functional with respect to time into the problem of solving the partial derivatives of the drift term function and the diffusion term function of SDEs of the stochastic process with respect to its independent variables.
@@
For solving the transformed problem of the $\mathcal{A}$ operator, it is required that the functional of the stochastic process is known and there is a strict requirement on the continuity of the functional (at least the second-order partial derivatives are required to be continuous). However, in many scenarios, the functional may not be known, and only the data for the inputs and outputs of that functional are available. In designing a Critic network for reinforcement learning, some activation functions such as Relu may not satisfy the continuity requirement, suggesting that the choice of the activation function may result in the functional of the value function that does not satisfy the conditions for the application of the $\mathcal{A}$ operator, which is a significant limitation for the design of the network. The $\mathcal{Y}$ operator proposed in this paper transforms this problem into a problem of solving the partial derivatives of the drift term function and the diffusion term function of SDEs obeyed by the random variable itself with respect to its independent variables. The continuity requirement of the pairwise value function is transformed into the continuity requirement of the drift term function and the diffusion term function of SDEs obeyed by the random variable itself.
The proposed $\mathcal{Y}$ operator allows more selectivity in the design of the Critic network when the SDEs obeyed by the system state satisfy the continuity condition. The activation functions of network such as the Relu function can be chosen. Moreover, the $\mathcal{Y}$ operator can be used as a reference tool in areas such as inverse reinforcement learning, offline reinforcement learning, to provide researchers with inspiring proposals. Therefore, the proposed $\mathcal{Y}$ operator effectively addresses the shortcomings articulated in Problem II.
@@ 8.1 The Proof of Proposition 1
By denoting the dimension of $X^{(1)}_{t}$ as $d_{x^{1}}$, the dimension of $X^{(2)}_{t}$ as $d_{x^{2}}$, and so on, and the dimension of $X^{(m)}_{t}$ as $d_{x^{m}}$, there are, in essence, $\sum_{i=1}^{m}d_{x^{i}}$ independent variables in the function $\Psi(X^{(1)}_{t},X^{(2)}_{t}\dots X^{(m)}_{t})$. Denoted $\sum_{i=1}^{m}d_{x^{i}}$ as $d_{M}$
According to the Taylor expansion formula for multivariate functions, (34) in Proposition 1 can be written in the following form:
@@
    Case1($i\neq k$): Since $dW^{(1)}_{t}$, $dW^{(2)}_{t} \dots$ $dW^{(m)}_{t}$ are independent of each other, it is clear that the following equation holds
@@ 8.2 The Proof of Proposition 3
According to Theorem 1, the (36) naturally holds.
And according to Kolmogorov forward equation,